### v1

# Beyond Exact Match: Semantically Reassessing Event Extraction by Large Language Models

Abstract
Event extraction has gained extensive research attention due to its broad range of applications. However, the current mainstream evaluation method for event extraction relies on token-level exact match, which misjudges numerous semantic-level correct cases. This reliance leads to a significant discrepancy between the evaluated performance of models under exact match criteria and their real performance. To address this problem, we propose RAEE, an automatic evaluation framework that accurately assesses event extraction results at semantic-level instead of token-level. Specifically, RAEE leverages Large Language Models (LLMs) as automatic evaluation agents, incorporating chain-of-thought prompting and an adaptive mechanism to achieve interpretable and adaptive evaluations for precision and recall of triggers and arguments. Extensive experimental results demonstrate that: (1) RAEE achieves a very high correlation with the human average; (2) after reassessing 14 models, including advanced LLMs, on 10 datasets, there is a significant performance gap between exact match and RAEE. The exact match evaluation significantly underestimates the performance of existing event extraction models, particularly underestimating the capabilities of LLMs; (3) fine-grained analysis under RAEE evaluation reveals insightful phenomena worth further exploration. The evaluation toolkit of our proposed RAEE will be publicly released.
Beyond Exact Match: Semantically Reassessing Event Extraction by Large Language Models
Anonymous ACL submission
1 Introduction
Event extraction (EE) aims to detect event instances as well as all of its participants and attributes by analyzing and identifying mentions of semantically defined entities and relationships within them Grishman (2015); Wei and Wang (2019). It is widely divided into two subtasks, Event Detection (ED) and Event Argument Extraction (EAE). While ED aims to identify all triggers that most explicitly convey the occurrence of an event, EAE extracts arguments with semantic roles corresponding to the trigger. Since EE benefits a broad range of applications Zhang et al. (2020); Han et al. (2021), it has gained extensive research attention in recent years Lin et al. (2020); Lu et al. (2022); Yan et al. (2023). Additionally, as large language models (LLMs) are increasingly used as agents Chen et al. (2024); Liu et al. (2023a), the need for accurate event extraction becomes more important, which helps in effectively tracking and labeling the actions of agents in real-world scenarios Park et al. (2023).
Figure 1: One case of EAE is misjudged by exact match (EM) evaluation but reassessed as correct by our proposed RAEE evaluation. With a given trigger (walk), EAE aims to extract corresponding arguments with semantic roles.
Currently, the most widely-used evaluation method for EE is token-level exact match (EM) [^1] For the sake of brevity, we will refer to exact match evaluation as EM in the following content., which assesses whether the predicted triggers and arguments are identical to the annotated ground-truth in benchmarks Hsu et al. (2022); Lu et al. (2022); Hsu et al. (2023a). However, this EM evaluation often misjudges a significant number of cases that are semantically correct or convey the same meaning as ground-truth but do not match it literally Li et al. (2021). For example, one case of EAE task, which extracts event arguments corresponding to a given trigger, is as shown in Figure 1. The predicted arguments, Path: on the Star Ferry terminal concourse and Direction: east, are not the same as the ground-truth and are evaluated as failure extractions under EM. However, it is obvious to assess them as semantically correct extractions. Also, co-reference instances are misjudged by EM, as shown in the example where Promenade and the past Kowloon - Canton Railway Terminus refer to the same thing. Consequently, models that may appear to perform poorly under EM actually have strong performance, indicating that the capabilities of existing EE models might be significantly underestimated. This issue is particularly pronounced in generative models, especially LLMs Huang et al. (2024), where the generations are often semantically correct.
To address this severe issue, we propose a reliable automatic evaluation framework designed to ReAssess Event Extraction models at semantic-level instead of token-level, named RAEE. Specifically, advanced LLMs are utilized to automatically assess the quality of EE by considering their semantic correctness within their context, such as the source document or sentence. Also, an adaptive prompting mechanism is designed to guide LLMs in generating interpretable and adaptive evaluation results for precision and recall of triggers and arguments, in a chain-of-thought approach Wei et al. (2023).
Meta-evaluation results on 8 models, including advanced LLMs, across 10 datasets demonstrate that RAEE achieves a very high correlation with human judgments and significantly outperforms EM. This indicates that our proposed RAEE is a more reliable EE evaluation framework. Additionally, we conduct a comprehensive reassessment for 14 models on 10 datasets using our proposed RAEE evaluation metrics. Extensive experiments reveal a significant performance gap between EM and RAEE, suggesting that existing baselines, particularly LLMs, have been greatly underestimated. Although LLMs still slightly underperform compared to smaller fine-tuned models under RAEE evaluation, the performance gap has narrowed considerably. Furthermore, the ranking order among these baselines has shifted significantly after reassessing under RAEE on most datasets. For example, on ED task, Tagprime-C Hsu et al. (2023a) and DEGREE Hsu et al. (2022) significantly outperform other baselines that perform better under EM. Finally, a fine-grained analysis of the correct and incorrect cases identified by RAEE uncovers intriguing and insightful phenomena worth future exploration. For example, while most spans are accurately matched, their event and role types are usually incorrect under RAEE evaluation. In summary, the contributions of our study are as follows:
  • RAEE: We propose a reliable evaluation framework for the event extraction task, leveraging advanced LLMs along with chain-of-thought prompting and an adaptive mechanism, demonstrating strong correlations with human judgments.
  • Comprehensive evaluation: We comprehensively reassess the performance of previous works and LLMs for the event extraction task using our proposed RAEE framework. Experimental results reveal that the capabilities of event extraction models are significantly underestimated, especially of generative models and LLMs.
  • Fine-grained analysis: We conduct a fine-grained analysis to uncover the correct and incorrect cases identified by RAEE, highlighting areas worth further exploration.
2 Related Works
2.1 Evaluations on Event Extraction
Current evaluations of EE typically use precision, recall, and f1-score metrics to evaluate the span and the type of extracted instances, and this approach is consistent across ED and EAE. The majority of research relies on EM evaluation Lin et al. (2020); Ma et al. (2022); Hsu et al. (2023b), though some studies employ language toolkits to assist with token-level evaluation Li et al. (2021). Additionally, some work modifies the existing evaluation frameworks by incorporating different averaging methods for calculating f1-score Zheng et al. (2021) or introducing additional metrics from downstream applications Peng et al. (2023a). The consistency and fairness of variable alignment with model outputs are also emphasized in recent studies Peng et al. (2023b); Huang et al. (2024). However, these researches uniformly adopt EM when determining span matches, overlooking instances where the span may not exactly match but is semantically correct. To address this oversight, we propose a semantic-level evaluation framework RAEE and provide a more comprehensive evaluation of model performance.
2.2 LLMs as Judge Models
There is an emerging trend in the research community to leverage large language models (LLMs) for automatically evaluating the quality of generations, thereby reducing the significant cost of human annotations Lan et al. (2024); Lin et al. (2024); Luo et al. (2023); Wang et al. (2023); Liu et al. (2023b); Fu et al. (2023). Extensive studies Fu et al. (2023); Wang et al. (2023) have shown that advanced LLMs, such as GPT-4, achieve very close correlations with human judgments across various generative NLP tasks, including dialogue generation Fu et al. (2023), text summarization Liu et al. (2023b) and structured generation Wang et al. (2023).
In this paper, to address the limitations of EM, we utilize advanced LLMs to mimic human annotators in automatically evaluating the quality of model-generated trigger words and arguments for EE task. Advanced LLMs achieve performance that is closely aligned with human judgments while incurring significantly lower costs Lan et al. (2020), making our proposed RAEE a scalable and reliable automatic evaluation method for EE task.
3 Automatic Reassessing Framework RAEE
To address the misjudgments caused by EM, we propose an automatic evaluation framework, RAEE, to reassess existing event extraction methods more accurately. In this section, we will introduce: (1) how RAEE differs from EM in Section 3.1; (2) how RAEE automatically achieves interpretable and adaptive evaluations in Section 3.2; (3) the reliability of RAEE in Section 3.3.
3.1 Semantic-level F1-score
Figure 2 shows the difference between RAEE and EM. In EM, True Positive (TP) cases are identical in the calculation of precision and recall. While in RAEE, we design the semantic-level F1-score by calculating precision and recall separately without counting TP as follows:
$\bf p=\frac{|P_{c}|}{|P|},\bf\quad r=\frac{|G_{r}|}{|G|},\quad\bf f1_{\rm RAEE}=\frac{2*p*r}{p+r}$
where $\bf P$ and $\bf G$ are predicted results and golden annotations. $\bf P_{c}$ and $\bf G_{r}$ are correct predicted results and recalled golden annotations, respectively.
Figure 2: Comparison of EM and RAEE evaluation using EAE as an example (consistent to ED). RAEE utilizes a LLM to assess event extraction results from both precision and recall perspectives. The source of the arrows indicates the basis or reason for judgment, with solid arrows representing the precision evaluation process and dashed arrows representing the recall evaluation process.
Since human evaluation is the most reliable golden standard Wang et al. (2023); Fu et al. (2023), the most accurate results for $\bf P_{c}$ and $\bf G_{r}$ should ideally be obtained through human annotation. However, due to the high cost, lengthy duration, and lack of reproducibility associated with human evaluation Lan et al. (2024, 2020), the advanced LLM, GPT-4o, is utilized as a reliable proxy for human evaluation in RAEE. Specifically, LLMs are equipped with an adaptive mechanism and chain-of-thought prompting strategies to evaluate the quality of event extraction results by considering the semantic correctness within context.
3.2 Adaptive Evaluation by LLMs
Human annotators might have slightly different subjective perspectives when evaluating, which could lead to differences in assessment Lan et al. (2024); Kim et al. (2023). Therefore, for both ED and EAE tasks, we manually identify several key criteria of evaluation bias, like evaluating pronouns in co-reference as correct extractions, and prompt LLMs to follow them to judge the quality of event extraction results step by step. More details about our human-annotated key criteria of evaluation bias and reassessment prompts can be found in Appendix C and Appendix D.1, respectively.
It should be noted that these key criteria are optional and adaptive, contributing to the adaptive mechanism in prompts. This allows researchers to tailor the evaluation criteria of RAEE to their specific needs in different scenarios. For example, some datasets provide separate annotations for co-references to assist in the annotation of event extraction Li et al. (2021) but some do not Ebner et al. (2020); Satyapanich et al. (2020). To achieve fair comparison in experiments, the evaluation criteria remain the same for all the evaluated models in our work.
Task - Agree. $\uparrow$ Spearman $\uparrow$
ED-p Human-Average 89.08 74.18
GPT4o-Human 76.89 49.09
      - w/o. Adaptive 78.99 47.25
EM-Human 26.05 ——
ED-r Human-Average 96.87 92.64
GPT4o-Human 87.17 74.72
      - w/o. Adaptive 84.19 69.75
EM-Human 71.05 ——
EAE-p Human-Average 93.61 85.47
GPT4o-Human 88.13 72.55
      - w/o. Adaptive 77.28 53.16
EM-Human 63.49 ——
EAE-r Human-Average 96.96 94.00
GPT4o-Human 86.10 75.27
      - w/o. Adaptive 86.42 76.08
EM-Human 72.55 ——
Table 1: Agreement percent and Spearman correlation with human average. -p and -r represent the evaluation of precision and recall, respectively. All Spearman correlation scores are associated with the $p$-value $<$ 0.05, indicating the results are statistically significant. Experimental results denote that LLM achieves a moderate and strong correlation with human judgments. Since EM fails to achieve basic correlation scores with human judgments, their Spearman correlation scores cannot be computed correctly.
3.3 RAEE Correlation with Human Evaluation
To prove the effectiveness of our proposed RAEE evaluation framework and adaptive evaluation, we conduct the meta-evaluation to analyze the correlation between RAEE and human judgments. Specifically, we invite 3 professional event extraction annotators to evaluate 200 sampled items of data for both ED and EAE tasks on precision and recall metrics. These annotators are given the same prompts as those given to LLM. With reassessing results from 3 humans and a LLM (GPT-4o), we calculate two widely-used inter-annotator agreement (IAA) metrics among humans and LLM: percent agreement Zheng et al. (2023); Liu et al. (2021) and Spearman correlation Lan et al. (2020, 2024); Fu et al. (2023). The experimental results are shown in Table 1.
3.3.1 Correlation on Recall Task
While LLM correlation with human average shows a noticeable drop (average 18.33 Spearman and 10.28% agreement percent), they still achieve a very strong correlation with human judgments on Agreement percent and Spearman correlation, i.e., over 86% agreement percent, which significantly outperforms EM with over 14.84% agreement percent. This also suggests that while the LLM can simulate human judgments and be regarded as an automatic evaluation agent, there is still room for further refinement.
3.3.2 Correlation on Precision Task
LLMs achieve a strong correlation with human judgments, over 76% agreement percent and 0.49 Spearman correlation of ED-p Lan et al. (2020); Fu et al. (2023). Furthermore, the average correlation scores between LLMs and human average for ED-p are significantly lower than that in the other tasks, indicating that semantic evaluation in this task is more challenging. Our investigation reveals two main reasons for this phenomenon: (1) Within the same domain, different datasets have personalized definitions for event types with similar names; (2) As for datasets across different domains, there are discrepancies in how events are defined as being triggered. When sampling data from multiple datasets, these subtle definitional differences can lead to confusion among evaluators when determining whether trigger predictions are accurate, leading to lower consistency in ED-p. In contrast, evaluating recall is simpler since it only requires determining whether gold annotations are retrieved, without judging the accuracy of the event type. Also, evaluating event arguments is more straightforward, as the role types for event arguments are clearly defined across datasets. This explains the relatively higher IAA observed in the three tasks other than ED-p.
3.3.3 Ablation of Adaptive Mechanism
Moreover, it can also be observed that the adaptive mechanism significantly improves the percent agreement and Spearman correlation between RAEE and human average in most cases, i.e., over 2.85% percent agreement and 26.34 Spearman correlation absolute improvements. This is because pre-defined criteria help reduce the randomness in the LLM evaluation.
Summary
While LLM slightly underperforms human average, it significantly outperforms EM and achieves a strong correlation with human judgments. It is important to note that RAEE is a trade-off solution by considering the evaluation cost and reliability, which could reliably assess the performance of event extraction results with affordable cost. For example, as shown in Table 2, the average time and financial cost of GPT-4o on each evaluated model and each dataset are much smaller than those of pure human annotation, proving the scalability and reliability of LLM judge models in our proposed RAEE framework.
Evaluation Cost Time (hours) Financial (\$)
RAEE + GPT-4o $\approx 0.23$ $\leq 5$
RAEE + Human $\approx 19.22$ $\approx 182.32$
Table 2: The average evaluation cost of human annotation and GPT-4o for each evaluated model on each dataset.
Model ACE05 Genia2011 CASIE PHEE
p r f1 p r f1 p r f1 p r f1
Token-level Evaluation under Exact Match
Tagprime-C 62.69 74.60 68.13 75.16 71.91 73.50 68.29 71.85 70.02 72.56 70.39 71.46
OneIE 70.94 70.69 70.82 75.52 72.30 73.87 69.68 72.32 70.98 74.86 67.10 70.77
Query&Extract 66.44 68.92 67.65 68.10 69.20 68.65 47.90 58.36 52.61 46.90 67.10 55.21
CEDAR 60.30 71.76 65.53 72.19 64.88 68.34 70.05 68.06 69.04 69.88 71.49 70.68
DEGREE 60.41 73.18 66.18 54.42 70.97 61.60 57.43 65.85 61.35 68.36 68.49 68.43
Semantic-level Evaluation under RAEE
Tagprime-C 85.82 81.71 83.71 94.15 88.39 91.18 92.65 93.53 93.09 87.77 78.51 82.88
OneIE 89.30 77.98 83.26 93.34 86.82 89.97 93.77 94.09 93.93 89.10 74.63 81.22
Query&Extract 87.16 76.38 81.41 91.88 85.90 88.78 81.75 91.88 86.52 66.83 75.52 70.91
CEDAR 79.10 78.15 78.63 92.05 82.22 86.86 93.83 90.46 92.11 84.21 78.81 81.42
DEGREE 82.70 81.53 82.11 88.95 87.84 88.39 92.71 90.54 91.61 86.07 80.20 83.03
Model Fewevent MEE-en M2E2 Average
p r f1 p r f1 p r f1 p r f1
Token-level Evaluation under Exact Match
Tagprime-C 63.62 66.00 64.79 79.27 82.18 80.70 46.35 62.07 53.07 66.85 71.28 68.81
OneIE 68.38 62.65 65.39 78.22 82.52 80.31 52.10 50.00 51.03 69.96 68.23 69.02
Query&Extract 67.42 60.68 63.88 77.22 82.18 79.62 47.70 65.52 55.21 60.24 67.42 63.26
CEDAR 44.18 62.26 51.68 81.71 79.37 80.52 45.54 52.87 48.94 63.41 67.24 64.96
DEGREE 60.33 59.78 60.05 78.06 78.68 78.37 43.46 59.20 50.12 60.35 68.02 63.73
Semantic-level Evaluation under RAEE
Tagprime-C 88.24 73.87 80.42 93.42 86.99 90.09 94.42 64.94 76.96 90.93 81.13 85.48
OneIE 90.03 69.85 78.67 93.21 86.65 89.81 96.41 53.45 68.77 92.17 77.64 83.66
Query&Extract 89.37 67.57 76.96 92.95 87.11 89.93 93.72 69.54 79.84 86.24 79.13 82.05
CEDAR 61.83 69.30 65.35 94.40 83.27 88.48 89.11 57.47 69.88 84.93 77.10 80.39
DEGREE 86.38 67.69 75.90 93.12 84.81 88.77 96.20 68.39 79.95 89.45 80.14 84.25
Table 3: Evaluation under EM versus reassessment under RAEE on ED task. The values in the Average column are derived from the mean of the values across 7 datasets. For each metric in each evaluation framework, we highlight the highest value using a dark shade, and the second-highest value using a light shade.
4 Re-Assessment Experiments
The reliability of RAEE with GPT-4o as the judging agent has been proven. In this section, we reassess the performance of existing event extraction models in numerous datasets across various domains by using our proposed RAEE evaluation framework.
4.1 Experimental Settings
Tasks and Metrics
ED and EAE event extraction tasks are evaluated respectively. Following Lin et al. (2020), we focus on trigger classification (TC) and argument classification (AC). The semantic-level f1-score for TC and AC is described in Section 3.1, using specialized prompts to assess both precision and recall.
Reassessed Models
We reassess the overall 8 state-of-the-art event extraction models with different frameworks (5 models for ED and EAE tasks). These models can be broadly classified into two categories: (1) Discriminative models that classify each token of text into predefined categories, including Tagprime-C Hsu et al. (2023a), OneIE Lin et al. (2020), Query&Extract Wang et al. (2022), CEDAR Li et al. (2023a); (2) Generative models that produce event-related outputs directly, including BartGen Li et al. (2021), PAIE Ma et al. (2022), DEGREE Hsu et al. (2022), AMPERE Hsu et al. (2023b). All models train and inference under the unified TextEE framework Huang et al. (2024), with implementation details shown in Appendix E.
Event Datasets
We use an overall 10 datasets to test the performance of models, covering diverse domains, including news of ACE05-en Doddington et al. (2004) and RAMS Ebner et al. (2020), general domain of FewEvent Deng et al. (2020) and GENEVA Parekh et al. (2023), Wikipedia of MEE-en Pouran Ben Veyseh et al. (2022) and WikiEvents Li et al. (2021), biomedical domain of Genia2011 Kim et al. (2011), multimedia of M2E2 Li et al. (2020), cybersecurity of CASIE Satyapanich et al. (2020), and Pharmacovigilance of PHEE Sun et al. (2022). The statistics of these datasets are provided in Appendix B.
Model ACE05 Genia2011 CASIE PHEE
p r f1 p r f1 p r f1 p r f1
Token-level Evaluation under Exact Match
Tagprime-C 75.23 70.05 72.55 83.46 68.79 75.42 66.39 66.07 66.23 71.69 45.84 55.92
BartGen 72.19 67.39 69.71 78.10 64.58 70.70 59.68 60.52 60.10 58.17 43.49 49.77
PAIE 74.14 70.29 72.16 81.07 67.84 73.87 65.13 62.90 64.00 72.99 74.34 73.66
DEGREE 74.30 67.39 70.68 70.08 57.95 63.44 55.29 51.32 53.23 61.42 47.63 53.65
AMPERE 69.87 64.13 66.88 71.45 56.89 63.35 59.08 50.21 54.29 60.94 44.54 51.46
Semantic-level Evaluation under RAEE
Tagprime-C 92.72 79.11 85.37 92.30 72.35 81.12 92.61 83.15 87.62 89.89 81.57 85.53
BartGen 91.34 79.59 85.06 91.26 73.79 81.60 86.21 80.74 83.39 84.16 79.26 81.64
PAIE 92.35 81.52 86.60 90.56 76.41 82.89 92.01 83.91 87.78 89.22 92.73 90.94
DEGREE 93.35 77.29 84.57 86.87 68.14 76.37 88.16 71.82 79.16 86.75 82.37 84.50
AMPERE 88.96 74.88 81.32 87.71 67.44 76.25 89.31 66.98 76.55 85.89 80.23 82.97
Model WikiEvents RAMS GENEVA Average
p r f1 p r f1 p r f1 p r f1
Token-level Evaluation under Exact Match
Tagprime-C 66.80 63.69 65.21 50.34 45.02 47.53 78.25 78.99 78.62 70.31 62.64 65.93
BartGen 70.24 63.69 66.80 48.70 43.37 45.88 64.20 63.69 63.94 64.47 58.10 60.99
PAIE 70.71 65.63 68.08 51.94 51.06 51.50 70.43 68.41 69.41 69.49 65.78 67.52
DEGREE 57.27 51.26 54.10 45.37 44.60 44.98 64.88 66.17 65.52 61.23 55.19 57.94
AMPERE 60.42 50.10 54.78 47.87 46.63 47.24 65.41 66.62 66.01 62.15 54.16 57.71
Semantic-level Evaluation under RAEE
Tagprime-C 93.32 76.36 83.99 81.31 70.14 75.31 93.65 92.75 93.20 90.83 79.35 84.59
BartGen 93.99 76.94 84.61 81.49 68.72 74.56 90.34 83.36 86.71 88.40 77.49 82.51
PAIE 92.74 79.84 85.81 81.23 74.89 77.93 93.25 88.77 90.96 90.19 82.58 86.13
DEGREE 88.48 70.74 78.62 80.22 72.87 76.37 91.43 86.14 88.70 87.89 75.62 81.18
AMPERE 93.61 67.64 78.53 81.70 72.87 77.03 91.56 86.64 89.03 88.39 73.81 80.24
Table 4: Evaluation under EM versus reassessment under RAEE on EAE task. The values in the Average column are derived from the mean of the values across 7 datasets. For each metric in each evaluation framework, we highlight the highest value using a dark shade, and the second-highest value using a light shade.
4.2 Main Results
Table 3 and 4 show EM evaluations and RAEE reassessment results for ED and EAE tasks, respectively. Notably, across all models, RAEE reassessment shows a significant improvement in metrics compared to EM evaluations. On average, the precision of all models increased by more than 20%, with recall improvement exceeding 10% for ED and 15% for EAE. This indicates that EM evaluation severely underestimates model performance by overlooking a substantial number of semantically correct predictions, while RAEE provides a more accurate evaluation of model performance across diverse datasets.
Task Model Category Average $\Delta$ f1
ED Discriminative Models 16.38
Generative Models 20.52
EAE Discriminative Models 18.66
Generative Models 21.47
Table 5: Average f1-score difference between EM and RAEE evaluation. It can be found that generative models are significantly more underestimated compared to discriminative models.
Moreover, as shown in Table 5, compared to discriminative models, the performance gap between RAEE and EM is significantly larger for generative models, indicating that EM introduces a greater bias for generative models.
Furthermore, the rank of models across different datasets changes considerably under RAEE. For instance, in the ED task, DEGREE ranks lower than OneIE and CEDAR on average under EM evaluation, while DEGREE outperforms these two models under RAEE. This suggests that EM evaluation distorts the true performance rank of models, leading to misconceptions about the model performance.
4.3 Fine-grained Reason Analysis
We further conduct a fine-grained analysis to uncover the main reasons behind the misjudgments by EM, as well as the primary failure modes of existing event extraction models under our proposed RAEE evaluation framework. Specifically, 200 sampled data items for both ED and EAE tasks are annotated through multiple human experts.
Reasons of EM Misjudgements
As for instances evaluated as incorrect under EM but reassessed as correct under RAEE, we summarize 4 main categories of reasons for ED and EAE tasks, through human annotations: (1) With CoreWord: Compared to the ground-truth, the predicted results contain the core words and only have missing or redundant words that do not affect the overall semantics; (2) Co-reference: There is a relationship between the predicted result and a certain ground-truth refer to the same entity or concept; (3) Correct: While no match is found in ground-truth, these instances are still reassessed as correct; (4) Other non-representative reasons.
Figure 3: Distribution of the reasons that misjudges by EM evaluation method on ED and EAE tasks.
As shown in Figure 3, it can be observed that the most misjudgments for EM evaluation in ED and EAE tasks are CoreWord and Correct. This indicates that current event extraction models tend to retain core words while expanding or omitting some less relevant words without changing the semantic meaning of the extracted results. Additionally, this also suggests that the annotations of existing datasets may not be comprehensive, as there are still a significant number of extracted results that are reasonable and correct but are not included in the ground-truth.
Failure Modes of Existing Models Under RAEE
As for instances evaluated as incorrect under both EM and RAEE, we also summarize 4 kinds of reasons for ED and EAE tasks, through human annotations: (1) Without CoreWord: Compared to the ground-truth, the predicted results have missing core words that convey the primary semantics; (2) WrongType: Even though there is a semantic-level match of span in the ground-truth, the predicted results have an incorrect type of event or role; (3) Incorrect: These instances are reassessed as incorrect and they are not included in the ground-truth; (4) Other non-representative reasons.
Figure 4: Distribution of failure modes under our proposed RAEE evaluation framework on ED and EAE tasks.
As shown in Figure 4, the most frequent failure modes of existing models on ED and EAE tasks are WrongType. This phenomenon indicates that although existing models could locate the spans correctly, their types are usually misclassified.This suggests that future works should focus on improving the capabilities of models in understanding and classifying the types of triggers and arguments.
Model RAEE EM
p r f1 f1
Tagprime-C 93.75 86.38 89.91 78.26
DEGREE 89.31 85.91 87.58 72.48
GPT-4o 79.08 53.05 63.50 32.76
        - 2 shots 77.66 48.35 59.60 33.17
        - 4 shots 78.77 49.76 60.99 28.23
        - 8 shots 74.41 61.50 67.34 37.79
GPT-3.5 55.69 45.07 49.82 13.94
Claude-3-Opus 76.31 40.84 53.21 13.69
Llama-3.1-70B 72.60 54.46 62.23 16.66
Qwen2-70B-Chat 61.53 55.39 58.30 14.47
Mixtral-8x7B 45.84 50.23 47.94 2.72
Table 6: LLMs versus Top-2 state-of-the-art smaller fine-tuned models on ED task under RAEE evaluation.
Model RAEE EM
p r f1 f1
PAIE 89.90 86.00 87.91 67.76
Tagprime-C 91.93 81.81 86.58 65.60
GPT-4o 83.24 73.19 77.89 15.89
        - 2 shots 82.49 69.56 75.48 16.36
        - 4 shots 81.75 69.02 74.85 16.38
        - 8 shots 80.02 69.92 74.63 15.57
GPT-3.5 62.87 64.13 63.49 10.26
Claude-3-Opus 78.82 67.39 72.66 20.22
Llama-3.1-70B 60.62 69.56 64.78 12.86
Qwen2-70B-Chat 79.95 68.11 73.56 16.99
Mixtral-8x7B 62.17 65.94 64.00 11.76
Table 7: LLMs versus Top-2 state-of-the-art smaller fine-tuned models on EAE task under RAEE evaluation.
5 Do LLMs Solve EE Task?
Under EM evaluation system, LLMs have a significant performance gap with state-of-the-art small fine-tuned models on EE task Huang et al. (2024). In this section, we also reassess the capabilities of LLMs. Specifically, 6 advanced open-source and closed-source LLMs are included: GPT-4o, GPT-3.5-turbo, Claude-3-Opus, Llama-3.1-70B-Chat, Qwen2-70B-Chat Yang et al. (2024) and Mixtral-8X7B-instruct Jiang et al. (2024). The well-selected prompts for inference on ED and EAE tasks are shown in Appendix D.2.
Figure 6 and 7 demonstrate results of LLMs and Top-2 fine-tuned small models under EM and RAEE evaluation on ED and EAE tasks, respectively. It can be found that LLMs generally underperform compared to fine-tuned models under EM evaluation, with significantly lower f1-scores on ED and EAE tasks. However, LLMs show remarkable improvements under RAEE evaluation, with f1-scores much closer to those of fine-tuned models, indicating that the event extraction capabilities of LLMs have been greatly underestimated. Notably, f1-score improvement of LLMs on EAE task, over 50%, is higher than that on ED task, over 30%. Our human annotations over hundreds of cases reveal that LLMs are more likely to generate out of pre-defined event type labels in ED task than role type labels in EAE task, as reflected in their relatively lower recall scores of ED.
Additionally, we also conduct few-shot experiments with the best-performing LLM, GPT-4o [^2] https://openai.com/index/hello-gpt-4o/. It is found that as the number of few-shot examples increases, the performance of the model fluctuates on ED task and declines on EAE task. Our human annotations reveal that these results are heavily influenced by the quality of selected shots. With data to be extracted and sampled from datasets in a diverse range of domains, it is more likely to mislead the model by providing a few shots from only a few specific domains. We will leave a more detailed analysis of few-shot prompting strategies in LLMs for our future work.
6 Conclusion and Future Works
In this paper, we propose a reliable automatic evaluation framework, RAEE, for assessing the event extraction task at the semantic level rather than the token level of exact match evaluation. Extensive experimental results demonstrate that GPT-4o achieves a strong correlation with human judgments within our RAEE framework, establishing RAEE as a reliable automatic evaluation framework for assessing the quality of event extraction results. Additionally, we reassess the performance of numerous prior works and advanced LLMs in the event extraction task using our proposed RAEE framework. Extensive experimental results indicate that the capabilities of prior works and LLMs have been significantly underestimated. Finally, the fine-grained reason analysis reveals several intriguing phenomena worth further exploration.
In the future, we will conduct a more comprehensive evaluation for event extraction tasks to reassess the performance of existing models. Besides, we will extend the evaluation of model performance with RAEE from closed domains to the open domain.
Appendix A Limitations
A.1 LLM Judgments v.s. Human Judgments
In this paper, advanced LLM GPT-4 is used to automatically judge the correctness of the predicted trigger words and arguments for the event extraction task. However, it should be noted that LLMs cannot always perfectly align with human judgments, which is the most reliable evaluation. For example, GPT-4o achieves comparable agreement and correlations with human average performance in ED and EAE annotations. Since the ED precision task is challenging, the Spearman correlation score on it is significantly lower than human performance. While we have made efforts to improve the relevance between LLM and human assessment through prompt design, this gap still exists. In the future, we would like to discover other strategies to further improve the reliability of LLM-as-a-Judge, such as multi-agent debate Chan et al. (2023).
A.2 Computational Cost of Evaluating One Model
Unlike existing EM evaluation, our proposed metrics rely on advanced LLMs to annotate the quality of predicted events, leading to much higher costs. The average cost of evaluating one model on each benchmark is between 2-5\$, which is smaller than established benchmarks like AlpacaEval Li et al. (2023b).
A.3 Insufficient Evaluation for LLMs
The number of the test set in existing event extract datasets contains thousands of samples, leading to a huge cost for LLMs to do inference experiments. Following previous works Huang et al. (2024), we sample a subset of the test set in existing datasets and prompt LLMs to handle the event extraction task. However, we wish to emphasize that the subset consists of diverse samples from 10 event extraction datasets. Therefore, the conclusions of our experimental results are reliable and robust.
A.4 Other Advanced LLMs as Evaluator
In this paper, the GPT-4 model is used to conduct the automatic evaluation in RAEE. Due to the budget limitation, we did not employ a diverse set of advanced LLMs for this task, such as Claude-3 and PaLM. Since our preliminary study reveals that GPT-4o is slightly better than other advanced LLMs, we mainly use it as the automatic judge model in our proposed RAEE framework.
Appendix B Datasets Statistics
We utilize 10 widely-used datasets as evaluation benchmarks in both EM and RAEE evaluation. Since these datasets come from different domains and vary in pre-defined ontologies and size, it is necessary to introduce basic statistics of these datasets, as shown in Table 8.
Dataset #Inst #ET #Evt #RT #Arg
ED and EAE tasks
ACE05-en 20920 33 5348 22 8097
Genia2011 1375 9 13537 10 11865
CASIE 1483 5 8469 26 22575
PHEE 4827 2 5019 16 25760
ED task only
FewEvent 12573 100 12573
MEE-en 13000 16 17257
M2E2 6013 8 1105
EAE task only
WikiEvents 565 50 3932 58 5501
RAMS 9647 139 9647 65 21206
GENEVA 3684 115 7505 220 12314
Table 8: Statistics of 10 datasets used in RAEE. #Inst, #ET, #Evt, #RT, and #Arg represent the number of data instances, event type, event instances, role types, and argument instances, respectively.
Appendix C Key Criteria in RAEE
Considering that the requirements for EE vary across different scenarios and applications, we introduce key criteria in RAEE to enable the adaptive mechanism. As shown in Tabel 9, these key criteria are based on factors that significantly influence the evaluation results in our human annotation meta-evaluation experiment. Furthermore, this adaptive mechanism is easily extendable, allowing for the addition of personalized judgment criteria to account for more factors in extraction.
Key Criteria Template Setting
If there is a more accurate predefined type for a predicted trigger/argument, then only the more accurate type can be reassessed as {} one correct
If a predict trigger/argument is more reasonable than golden annotations, it should be reassessed as {} one correct
When there is co-reference, pronouns can be reassessed as {} one correct
If the core word is correct but some modifiers are missing, the predicted trigger/argument should be reassessed as {} one correct
A predicted trigger is reassessed as {} one when it triggers an event that does not really occur incorrect
A golden trigger/argument is reassessed as {} one if there is no correspondence in predicted triggers/arguments, even if there is a predicted trigger/argument with a more accurate or reasonable type not recalled
When there is co-reference, pronouns can be reassessed as {} one recalled
If the core word is recalled but some modifiers are missing, the golden trigger/argument should be reassessed as {} one recalled
Table 9: Key criteria templates and their settings used in RAEE.
Appendix D Prompts
D.1 Reassessment Prompts in RAEE
Table 10, 11, 13 and 15 illustrate the prompts we use for reassessment in RAEE.
D.2 Inference Prompts for LLMs Extracting Events
Table 12 and 14 illustrate the prompts we use for testing the event extraction ability of LLMs in ED and EAE tasks, respectively.
Prompt Used for Reassessing on ED Task and Precision Metric
Instruction You are a reassessor designed to reassess predicted event triggers from an event detection model. Task Description: Refer to a piece of context and golden triggers, reassess predicted event triggers. A trigger is the key word in the sentence that most explicitly conveys the occurrence of the event. A predicted trigger can be reassessed as correct one if it conveys the occurrence of a particular interest kind of event, even if it does not exactly match any of golden triggers; otherwise, it is reassessed as incorrect one.
Key Criteria There are several rules to follow when reassessing: (1) If there is a more accurate predefined type for a predicted trigger, then only the more accurate type can be reassessed as correct one; (2) If a predict trigger is more reasonable than golden annotations, it should be reassessed as correct one; (3) When there is co-reference, pronouns can be reassessed as correct one; (4) If the core word is correct but some modifiers are missing, the predicted trigger should be reassessed as correct one; (5) A predicted trigger is reassessed as incorrect one when it triggers an event that does not really occur.
Requirements When answering, provide your thought process at first, including analyzing the context, explaining how golden triggers indicate events with a specific type, and reassessing predicted triggers one by one based on your thought. After that, answer final reassessment results.
Please answer in the following python dict format: {’Thought Process’: {’Context Analysis’: str, ’Gold Triggers’: [your explanation for gold trigger 0, your explanation for gold trigger 1, …], ’Predicted Triggers’: [your reassessment thought for predicted trigger 0, your reassessment thought for predicted trigger 1, …]}, ’Final Reassessment Results’: [1, 0, …]}.
Data Events of interest include:
Contact.Meet. The event is related to contact meet.
Business.Start-Org. The event is related to business start organization.
Context:
Professor Massimo Pigliucci , a [t.Pred.1] [t.Pred.0] professor [/t.Pred.0] [/t.Pred.1] of ecology and evolution at the State University of New York at Stony Brook , has criticized ICR for professing to present the same science as that taught in secular universities while at the same time requiring students and faculty to sign a statement of faith to ICR ’s fundamentalist religious mission , most notably in affirming conformity in all its [t.Gold.0] work [/t.Gold.0] to Biblical doctrine .
The position of each following trigger is labeled in the context between a pair of [t] and [/t].
Trigger Gold.0: work # event type: Business.Employment-Tenure
Trigger Pred.0: professor # event type: Business.Employment-Tenure
Trigger Pred.1: professor # event type: Education.Education
Table 10: Prompt used for reassessing on ED task and precision metric. An example of data is illustrated.
Prompt Used for Reassessing on ED Task and Recall Metric
Instruction You are a reassessor designed to reassess the recall result of an event detection model. Task Description: According to a piece of context, predicted triggers and golden triggers, reassess whether each golden trigger is recalled. A trigger is the key word in the sentence that most explicitly conveys the occurrence of the event. A golden trigger can be reassessed as recalled one if it can correspond to one or more predicted triggers that convey the occurrence of an event with the same interest type, even if it does not exactly match any of predicted triggers; otherwise, it is reassessed as not recalled one.
Key Criteria There are several rules to follow when reassessing: (1) A golden trigger is reassessed as not recalled one if there is no correspondence in predicted triggers, even if there is a predicted trigger with more accurate or reasonable type; (2) When there is co-reference, pronouns can be reassessed as recalled one; (3) If the core word is recalled but some modifiers are missing, the golden trigger should be reassessed as recalled one.
Requirements When answering, provide your thought process at first, including analyzing the context, explaining how golden triggers indicate events with a specific type, finding correspondence with predicted triggers for each golden trigger or explaining why any correspondence cannot be found, and reassessing golden triggers one by one based on your thought. After that, answer final reassessment results.
Please answer in the following python dict format: {’Thought Process’: {’Context Analysis’: str, ’Gold Triggers’: [your finding or explanation for gold trigger 0, your finding or explanation for gold trigger 1, …]}, ’Final Reassessment Results’: [1 , 0 , …]}.
Data Events of interest include:
Business.Start-Subsidiary. The event is related to business start subsidiary.
Context:
A subsidiary company or daughter company is a company that is [t.Gold.0] owned or [t.Pred.0] controlled [/t.Pred.0] by [/t.Gold.0] another company , which is called the parent company , parent , or holding company .
The position of each following trigger is labeled in the context between a pair of [t] and [/t].
Trigger Pred.0: controlled # event type: Business.Start-Subsidiary
Trigger Gold.0: owned or controlled by # event type: Business.Start-Subsidiary
Table 11: Prompt used for reassessing on ED task and recall metric. An example of data is illustrated.
Prompt Used for Event Detection
Instruction You are an event extractor designed to check for the presence of a specific event in a piece of context and to locate the corresponding event trigger. Task Description: Identify all triggers related to the event of interest in the context. A trigger is the key word in the context that most explicitly conveys the occurrence of the event.
Requirements When extracting, analyzing the context at first. After that, answer extraction results. You need to extract the span of a extracted trigger as well as its corresponding event type. Note that there may be zero to plural triggers in the context. Please answer in the following python dict format: {’Context Analysis’: str, ’Extraction Results’: [{’Trigger Span’: a span in the context, ’Event Type’: a specific event of interest}, …]}.
Data Types of events that may occur include:
Adverse_event. The event is related to adverse event.
Potential_therapeutic_event. The event is related to potential therapeutic event.
Context:
The case histories are presented of two patients who developed lung disease associated with the use of nitrofurantoin with histological features of bronchiolitis obliterans organising pneumonia ( BOOP ) , a rare but recognised form of drug induced injury .
Table 12: Prompt used for testing the event extraction ability of LLMs in ED task. An example of data is illustrated.
Prompt Used for Reassessing on EAE Task and Precision Metric
Instruction You are a reassessor designed to reassess predicted event arguments from an event argument extraction model. Task Description: Refer to a piece of context and golden arguments, reassess predicted event arguments. Arguments have the semantic role corresponding to the given event trigger by the word span between [t] and [/t]. A predicted argument can be reassessed as correct one if it have a particular interest type of role corresponding to the given event trigger, even if it does not exactly match any of golden arguments; otherwise, it is reassessed as incorrect one.
Key Criteria There are several rules to follow when reassessing: (1) If there is a more accurate predefined role type for a predicted argument, then only the more accurate role type can be reassessed as correct one; (2) If a predict argument is more reasonable than golden annotations, it should be reassessed as correct one; (3) When there is co-reference, pronouns can be reassessed as correct one; (4) If the core word is correct but some modifiers are missing, the predicted argument should be reassessed as correct one.
Requirements When answering, provide your thought process at first, including analyzing the context, explaining roles of golden arguments in the event of interest, and reassessing predicted arguments one by one based on your thought. After that, answer final reassessment results.
Please answer in the following python dict format: {’Thought Process’: {’Context Analysis’: str, ’Gold Arguments’: [your explanation for gold argument 0, your explanation for gold argument 1, …], ’Predicted Arguments’: [your reassessment thought for predicted argument 0, your reassessment thought for predicted argument 1, …]}, ’Final Reassessment Results’: [1, 0, …]}.
Data The event of interest is Justice:Trial-Hearing. The event is related to a trial or hearing for someone.
Argument roles of interest: [’Defendant’, ’Prosecutor’, ’Place’, ’Adjudicator’]
Context:
She was not [t] tried [/t] in the deaths of the other two .
Gold argument 0: She # role type: Defendant
Pred argument 0: deaths # role type: Defendant
Table 13: Prompt used for reassessing on EAE task and precision metric. An example of data is illustrated.
Prompt Used for Event Argument Extraction
Instruction You are an argument extractor designed to check for the presence of arguments regarding specific roles for an event in a piece of context. Task Description: Identify all event arguments related to roles of interest in the context. These arguments should have the semantic role corresponding to the given event trigger by the word span between [t] and [/t].
Requirements When extracting, analyzing the context at first. After that, answer extraction results. Note that there may be zero to plural arguments for each role. Please answer in the following python dict format: {’Context Analysis’: str, ’Extraction Results’: {’role1’: [str(the argument span), …(The length of this list depends on the number of arguments that you extract for this kind of role)], ’role2’: [str, …], …}}.
Data The event of interest is justice.initiatejudicialprocess.trialhearing. The event is related to justice initiative judicial process trial hearing.
Roles of interest include: prosecutor, defendant, judgecourt, crime, place.
Context:
But no longer . By the end of 2015 the establishment was expressing considerable dismay and desperation over its inability to do so , as the Republican base and its choices fell out of control . Republican elected officials and contenders for the next presidential election expressed open contempt for the Paris deliberations , refusing to even attend the proceedings . The three candidates who led in the polls at the time Donald Trump , Ted Cruz , and Ben Carson adopted the stand of the largely evangelical base : humans have no impact on global warming , if it is happening at all . The other candidates reject government action to deal with the matter .
Table 14: Prompt used for testing the event extraction ability of LLMs in EAE task. An example of data is illustrated.
Appendix E Implementation Details
To avoid inconsistency in evaluation caused by factors out of model frameworks such as data preprocessing or Python package versions, we follow TextEE Huang et al. (2024), utilize its standardized data splits and unified input-output framework for evaluating models, and adhere to its default training configs, which can be found in the corresponding GitHub repository. All discriminative models are based on the RoBERTa-large Liu et al. (2019) pre-trained model, and generative models are built using BART-large Lewis et al. (2019).
Prompt Used for Reassessing on EAE Task and Recall Metric
Instruction You are a reassessor designed to reassess the recall result of an event argument extraction model. Task Description: According to a piece of context, predicted arguments and golden arguments, reassess whether each golden argument is recalled. Arguments have the semantic role corresponding to the given event trigger by the word span between [t] and [/t]. A golden argument can be reassessed as recalled one if it can correspond to one or more predicted arguments that have the same interest type of role corresponding to the given event trigger, even if it does not exactly match any of predicted arguments; otherwise, it is reassessed as not recalled one.
Key Criteria There are several rules to follow when reassessing: (1) A golden argument is reassessed as not recalled one if there is no correspondence in predicted arguments, even if there is a predicted argument with more accurate or reasonable role type; (2) When there is co-reference, pronouns can be reassessed as recalled one; (3) If the core word is recalled but some modifiers are missing, the golden argument should be reassessed as recalled one.
Requirements When reassessing, provide your thought process at first, including analyzing the context, explaining roles of golden arguments in the event of interest, finding correspondence with predicted arguments for each golden argument or explaining why any correspondence cannot be found, and reassessing golden arguments one by one based on your thought. After that, answer final reassessment results.
Please answer in the following python dict format: {’Thought Process’: {’Context Analysis’: str, ’Gold Arguments’: [your finding or explanation for gold argument 0, your finding or explanation for gold argument 1, …]}, ’Final Reassessment Results’: [1, 0, …]}.
Data The event of interest is Life:Die. The event is related to life and someone died.
Argument roles of interest: [’Agent’, ’Victim’, ’Instrument’, ’Place’]
Context:
There ’s nothing very pretty about the training that you take to prepare you for combat , because it is to [t] kill [/t] people .
Pred argument 0: people # role type: Victim
Gold argument 0: you # role type: Agent
Table 15: Prompt used for reassessing on EAE task and recall metric. An example of data is illustrated.

### v2

@@
 This reliance leads to a significant discrepancy between the evaluated performance of models under exact match criteria and their real performance.
-To address this problem, we propose RAEE, an automatic evaluation framework that accurately assesses event extraction results at semantic-level instead of token-level.
-Specifically, RAEE leverages Large Language Models (LLMs) as automatic evaluation agents, incorporating chain-of-thought prompting and an adaptive mechanism to achieve interpretable and adaptive evaluations for precision and recall of triggers and arguments.
-Extensive experimental results demonstrate that: (1) RAEE achieves a very high correlation with the human average; (2) after reassessing 14 models, including advanced LLMs, on 10 datasets, there is a significant performance gap between exact match and RAEE.
-The exact match evaluation significantly underestimates the performance of existing event extraction models, particularly underestimating the capabilities of LLMs; (3) fine-grained analysis under RAEE evaluation reveals insightful phenomena worth further exploration.
-The evaluation toolkit of our proposed RAEE will be publicly released.
-Beyond Exact Match: Semantically Reassessing Event Extraction by Large Language Models
-Anonymous ACL submission
+To address this problem, we propose a reliable and semantic evaluation framework for event extraction, named RAEE, which accurately assesses extraction results at semantic-level instead of token-level.
+Specifically, RAEE leverages large language models (LLMs) as evaluation agents, incorporating an adaptive mechanism to achieve adaptive evaluations for precision and recall of triggers and arguments.
+Extensive experiments demonstrate that: (1) RAEE achieves a very strong correlation with human judgments; (2) after reassessing 14 models, including advanced LLMs, on 10 datasets, there is a significant performance gap between exact match and RAEE.
+The exact match evaluation significantly underestimates the performance of existing event extraction models, and in particular underestimates the capabilities of LLMs; (3) fine-grained analysis under RAEE evaluation reveals insightful phenomena worth further exploration.
+The evaluation toolkit of our proposed RAEE is publicly released [^1] https://github.com/Lyfralston/RAEE.
+Beyond Exact Match: Semantically Reassessing Event Extraction
+by Large Language Models
+Yi-Fan Lu1, Xian-Ling Mao1, Tian Lan1, Heyan Huang1, Chen Xu1, Xiaoyan Gao2 1Beijing Institute of Technology, 2Beijing University of Technology yifanlu@bit.edu.cn, maoxl@bit.edu.cn, lantiangmftby@gmail.com, chenxu05037@gmail.com, hhy63@bit.edu.cn, gaoxiaoyan@bjut.edu.cn
 1 Introduction
-Event extraction (EE) aims to detect event instances as well as all of its participants and attributes by analyzing and identifying mentions of semantically defined entities and relationships within them Grishman (2015); Wei and Wang (2019).
-It is widely divided into two subtasks, Event Detection (ED) and Event Argument Extraction (EAE).
-While ED aims to identify all triggers that most explicitly convey the occurrence of an event, EAE extracts arguments with semantic roles corresponding to the trigger.
+Figure 1: One case of EAE is misjudged by exact match (EM) evaluation but reassessed as correct by our proposed RAEE evaluation.
+With a given trigger (walk), EAE aims to extract its arguments with semantic roles.
+Event extraction (EE) aims to detect event instances as well as all of their participants and attributes by analyzing and identifying mentions of semantically defined entities and relationships within them Grishman (2015); Wei and Wang (2019).
+It is typically divided into two subtasks, Event Detection (ED), which focuses on identifying triggers that indicate event occurrence, and Event Argument Extraction (EAE), which extracts arguments with semantic roles related to the trigger.
 Since EE benefits a broad range of applications Zhang et al.
 (2020); Han et al.
-(2021), it has gained extensive research attention in recent years Lin et al.
-(2020); Lu et al.
+(2021), it has gained extensive research attention Lu et al.
 (2022); Yan et al.
 (2023).
-Additionally, as large language models (LLMs) are increasingly used as agents Chen et al.
-(2024); Liu et al.
-(2023a), the need for accurate event extraction becomes more important, which helps in effectively tracking and labeling the actions of agents in real-world scenarios Park et al.
+With the increasing use of large language models (LLMs) as agents Chen et al.
+(2024b); Liu et al.
+(2023a), the requirement for accurate event extraction becomes more important, which helps in effectively tracking and labeling the actions of agents in real-world scenarios Park et al.
 (2023).
-Figure 1: One case of EAE is misjudged by exact match (EM) evaluation but reassessed as correct by our proposed RAEE evaluation.
-With a given trigger (walk), EAE aims to extract corresponding arguments with semantic roles.
 Currently, the most widely-used evaluation method for EE is token-level exact match (EM) [^2] For the sake of brevity, we will refer to exact match evaluation as EM in the following content., which assesses whether the predicted triggers and arguments are identical to the annotated ground-truth in benchmarks Hsu et al.
@@
 (2023a).
-However, this EM evaluation often misjudges a significant number of cases that are semantically correct or convey the same meaning as ground-truth but do not match it literally Li et al.
+However, EM often misjudges a significant number of cases that are semantically correct or convey the same meaning as ground-truth but cannot match it literally Li et al.
 (2021).
-For example, one case of EAE task, which extracts event arguments corresponding to a given trigger, is as shown in Figure 1.
-The predicted arguments, Path: on the Star Ferry terminal concourse and Direction: east, are not the same as the ground-truth and are evaluated as failure extractions under EM.
-However, it is obvious to assess them as semantically correct extractions.
-Also, co-reference instances are misjudged by EM, as shown in the example where Promenade and the past Kowloon - Canton Railway Terminus refer to the same thing.
-Consequently, models that may appear to perform poorly under EM actually have strong performance, indicating that the capabilities of existing EE models might be significantly underestimated.
-This issue is particularly pronounced in generative models, especially LLMs Huang et al.
-(2024), where the generations are often semantically correct.
-To address this severe issue, we propose a reliable automatic evaluation framework designed to ReAssess Event Extraction models at semantic-level instead of token-level, named RAEE.
-Specifically, advanced LLMs are utilized to automatically assess the quality of EE by considering their semantic correctness within their context, such as the source document or sentence.
-Also, an adaptive prompting mechanism is designed to guide LLMs in generating interpretable and adaptive evaluation results for precision and recall of triggers and arguments, in a chain-of-thought approach Wei et al.
-(2023).
-Meta-evaluation results on 8 models, including advanced LLMs, across 10 datasets demonstrate that RAEE achieves a very high correlation with human judgments and significantly outperforms EM.
-This indicates that our proposed RAEE is a more reliable EE evaluation framework.
-Additionally, we conduct a comprehensive reassessment for 14 models on 10 datasets using our proposed RAEE evaluation metrics.
-Extensive experiments reveal a significant performance gap between EM and RAEE, suggesting that existing baselines, particularly LLMs, have been greatly underestimated.
-Although LLMs still slightly underperform compared to smaller fine-tuned models under RAEE evaluation, the performance gap has narrowed considerably.
-Furthermore, the ranking order among these baselines has shifted significantly after reassessing under RAEE on most datasets.
-For example, on ED task, Tagprime-C Hsu et al.
-(2023a) and DEGREE Hsu et al.
-(2022) significantly outperform other baselines that perform better under EM.
-Finally, a fine-grained analysis of the correct and incorrect cases identified by RAEE uncovers intriguing and insightful phenomena worth future exploration.
-For example, while most spans are accurately matched, their event and role types are usually incorrect under RAEE evaluation.
-In summary, the contributions of our study are as follows:
-  • RAEE: We propose a reliable evaluation framework for the event extraction task, leveraging advanced LLMs along with chain-of-thought prompting and an adaptive mechanism, demonstrating strong correlations with human judgments.
-  • Comprehensive evaluation: We comprehensively reassess the performance of previous works and LLMs for the event extraction task using our proposed RAEE framework.
+For example, one case of EAE task, which extracts event arguments corresponding to a given trigger, is illustrated in Figure 1.
+The predicted arguments, Path: on the Star Ferry terminal concourse and Direction: east, are not the same as ground-truth and are therefore evaluated as failed extractions under EM.
+However, considering the context, both predictions can be regarded as semantically correct extractions.
+Also, co-reference instances are misjudged by EM, as shown in the example where Promenade and the past Kowloon - Canton Railway Terminus refer to the same place.
+Consequently, models that may appear to perform poorly under EM actually have strong performance, indicating that the capabilities of existing EE models might be greatly underestimated.
+This issue is particularly prominent in generative models, especially LLMs Huang et al.
+(2024), where the generations often cannot exactly match ground-truth.
+To address this severe issue, we propose a reliable and semantic evaluation framework designed to ReAssess Event Extraction models at semantic-level instead of token-level, named RAEE.
+Specifically, RAEE utilizes advanced LLMs as evaluation agents to achieve automatic assessment.
+By considering the semantics of extractions in the context, the LLM performs binary judgments on the extractions from the perspectives of precision and recall, determining whether they are corrected and whether they are successfully recalled, respectively.
+To ensure the reliability of the evaluation, we introduce an adaptive prompting mechanism to enhance LLMs’ evaluation consistency with human evaluation.
+Experiments begin with a meta-evaluation designed to validate evaluation reliability of RAEE.
+The results demonstrate that (1) RAEE achieves a strong correlation with human judgments, (2) RAEE significantly outperforms EM and other soft matching evaluation methods, and (3) the proposed adaptive mechanism enhances the reliability of RAEE.
+Then, we apply RAEE to conduct a comprehensive reassessment experiment for 14 EE models on 10 widely-used datasets.
+Extensive experiments reveal that (1) evaluated models exhibit a significant performance gap under EM and RAEE, indicating that existing EE models, particularly generative models, are greatly underestimated, and (2) the ranking order among these models shifts after reassessing under RAEE on most datasets, further highlighting the importance of RAEE.
+To further explore the reasons behind misjudgments evaluated by EM and failure modes evaluated by RAEE, a fine-grained analysis experiment uncovers insightful phenomena worth future exploration.
+For example, while most spans are accurately matched, their types are still incorrect under RAEE evaluation.
+Finally, we reassess LLMs’ event extraction ability and find that LLMs are extremely underestimated by EM, further emphasizing the urgency of incorporating semantic evaluation in the era of LLMs.
+In summary, our contributions are as follows:
+  • RAEE: We propose a reliable and semantic evaluation framework to assess event extraction at semantic-level by leveraging advanced LLMs as evaluation agents with an adaptive prompting mechanism.
+RAEE shows strong correlations with human judgments.
+  • Comprehensive evaluation: We apply our proposed RAEE to comprehensively reassess the performance of previous works and LLMs for event extraction.
 Experimental results reveal that the capabilities of event extraction models are significantly underestimated, especially of generative models and LLMs.
-  • Fine-grained analysis: We conduct a fine-grained analysis to uncover the correct and incorrect cases identified by RAEE, highlighting areas worth further exploration.
+  • Fine-grained analysis: We conduct a fine-grained analysis to uncover the reasons behind misjudgments evaluated by EM and failure modes evaluated by RAEE, highlighting areas worth further exploration.
+Figure 2: Evaluation process of RAEE, using the precision of EAE as an example (consistent with ED).
 2 Related Works
@@
 (2022); Hsu et al.
-(2023b), though some studies employ language toolkits to assist with token-level evaluation Li et al.
+(2023b), and some studies employ language toolkits to apply head noun evaluation Li et al.
 (2021).
-Additionally, some work modifies the existing evaluation frameworks by incorporating different averaging methods for calculating f1-score Zheng et al.
+Additionally, some works modify the existing evaluation frameworks by incorporating different averaging methods for calculating f1-score Zheng et al.
 (2021) or introducing additional metrics from downstream applications Peng et al.
@@
 (2024).
-However, these researches uniformly adopt EM when determining span matches, overlooking instances where the span may not exactly match but is semantically correct.
-To address this oversight, we propose a semantic-level evaluation framework RAEE and provide a more comprehensive evaluation of model performance.
+However, these works uniformly adopt EM when determining span matches, and overlook instances where the semantically correct span cannot exactly match ground-truth.
+To address this oversight, we propose a semantic-level evaluation framework and provide a more reliable evaluation of model performances.
 2.2 LLMs as Judge Models
-There is an emerging trend in the research community to leverage large language models (LLMs) for automatically evaluating the quality of generations, thereby reducing the significant cost of human annotations Lan et al.
+There is an emerging trend in the research community to leverage LLMs for automatically evaluating the quality of generations, thereby reducing the significant cost of human annotations Lan et al.
 (2024); Lin et al.
 (2024); Luo et al.
+(2023); Fu et al.
 (2023); Wang et al.
-(2023); Liu et al.
-(2023b); Fu et al.
 (2023).
@@
 (2023).
-In this paper, to address the limitations of EM, we utilize advanced LLMs to mimic human annotators in automatically evaluating the quality of model-generated trigger words and arguments for EE task.
-Advanced LLMs achieve performance that is closely aligned with human judgments while incurring significantly lower costs Lan et al.
-(2020), making our proposed RAEE a scalable and reliable automatic evaluation method for EE task.
-3 Automatic Reassessing Framework RAEE
-To address the misjudgments caused by EM, we propose an automatic evaluation framework, RAEE, to reassess existing event extraction methods more accurately.
-In this section, we will introduce: (1) how RAEE differs from EM in Section 3.1; (2) how RAEE automatically achieves interpretable and adaptive evaluations in Section 3.2; (3) the reliability of RAEE in Section 3.3.
+In this paper, we utilize advanced LLMs to emulate human annotators in automatically evaluating EE.
+Advanced LLMs closely align with human judgments at a fraction of the cost Lan et al.
+(2020), making our RAEE a scalable and reliable automatic evaluation method for EE.
+3 Semantic Reassessing Framework
+To address the misjudgments caused by EM, we propose a semantic evaluation framework, named RAEE.
+In this section, we will introduce: (1) how semantic-level F1-score is calculated in RAEE in Section 3.1; (2) the adaptive mechanism of RAEE in Section 3.2; (3) the evaluation reliability of RAEE in Section 3.3.
 3.1 Semantic-level F1-score
-Figure 2 shows the difference between RAEE and EM.
-In EM, True Positive (TP) cases are identical in the calculation of precision and recall.
-While in RAEE, we design the semantic-level F1-score by calculating precision and recall separately without counting TP as follows:
-$\bf p=\frac{|P_{c}|}{|P|},\bf\quad r=\frac{|G_{r}|}{|G|},\quad\bf f1_{\rm RAEE}=\frac{2*p*r}{p+r}$
-where $\bf P$ and $\bf G$ are predicted results and golden annotations. $\bf P_{c}$ and $\bf G_{r}$ are correct predicted results and recalled golden annotations, respectively.
-Figure 2: Comparison of EM and RAEE evaluation using EAE as an example (consistent to ED).
-RAEE utilizes a LLM to assess event extraction results from both precision and recall perspectives.
-The source of the arrows indicates the basis or reason for judgment, with solid arrows representing the precision evaluation process and dashed arrows representing the recall evaluation process.
-Since human evaluation is the most reliable golden standard Wang et al.
-(2023); Fu et al.
-(2023), the most accurate results for $\bf P_{c}$ and $\bf G_{r}$ should ideally be obtained through human annotation.
-However, due to the high cost, lengthy duration, and lack of reproducibility associated with human evaluation Lan et al.
-(2024, 2020), the advanced LLM, GPT-4o, is utilized as a reliable proxy for human evaluation in RAEE.
-Specifically, LLMs are equipped with an adaptive mechanism and chain-of-thought prompting strategies to evaluate the quality of event extraction results by considering the semantic correctness within context.
-3.2 Adaptive Evaluation by LLMs
-Human annotators might have slightly different subjective perspectives when evaluating, which could lead to differences in assessment Lan et al.
+Assume there are a piece of context $\bf C$, predicted results from an EE model ${\bf P}={\rm Model}(\bf C)$, and gold annotations from human ${\bf G}={\rm Human}(\bf C)$.
+RAEE prompts an evaluation agent LLM to consider the context during evaluation, assessing predictions and annotations separately when calculating precision and recall.
+Specifically, RAEE requires the LLM to identify a correct prediction set $\bf P_{c}$ and a recalled annotation set $\bf G_{r}$:
+${\bf P_{c}}={\rm LLM}({\bf P|C,G}),\quad{\bf G_{r}}={\rm LLM}({\bf G|C,P}).$
+Then, semantic-level f1-score can be computed as follows:
+$\bf p=\frac{|P_{c}|}{|P|},\bf\quad r=\frac{|G_{r}|}{|G|},\quad\bf f1_{\rm RAEE}=\frac{2*p*r}{p+r}.$
+Figure 2 presents an example to evaluate predictions ${\bf P}$ in EAE.
+The LLM is required to first analyze the context and then evaluate each prediction with a reason to get correct predictions $\bf P_{c}$.
+Eval. Methods ED precision ED recall EAE precision EAE recall
+Per. Sp. Per. Sp. Per. Sp. Per. Sp.
+Human Average 89.08 74.19 96.87 92.65 93.61 85.48 96.96 94.00
+Deepseek-r1 77.451.65 53.192.57 86.850.22 72.310.54 79.720.42 55.170.78 93.490.53 87.021.08
+      - w/o. Adaptive 72.600.93 48.211.16 85.240.20 71.130.33 76.980.51 54.241.00 91.750.41 83.470.82
+GPT-4o-240513 78.290.50 50.492.31 86.850.80 71.701.74 84.630.69 64.932.27 83.710.55 68.810.54
+      - w/o. Adaptive 75.301.41 38.333.80 83.721.23 69.331.92 80.561.32 57.732.63 82.562.61 66.885.00
+Gemini-1.5-pro 72.830.89 37.932.26 51.490.77 34.600.73 81.350.61 55.031.48 77.160.22 61.810.44
+      - w/o. Adaptive 71.941.06 28.143.12 48.511.77 31.192.21 79.570.89 49.622.59 68.551.42 49.742.37
+Deepseek-v3 66.391.59 28.613.38 71.472.51 54.033.13 77.592.39 52.784.41 76.330.54 57.061.55
+      - w/o. Adaptive 66.711.49 31.071.91 61.872.95 44.043.37 72.830.28 48.771.71 72.101.93 49.872.52
+Claude-3.5-sonnet 63.631.09 33.982.96 71.620.92 51.771.30 75.610.57 48.651.31 71.881.01 50.751.29
+      - w/o. Adaptive 61.480.52 34.831.47 64.323.04 43.494.33 68.910.19 44.411.03 68.621.43 43.953.09
+Similarity 50.70 12.62 84.82 71.40 63.81 16.76 75.79 54.44
+Head Noun 31.93 11.80 75.74 29.97 40.98 20.07 64.39 35.98
+Exact Match 28.57 —— 72.46 —— 31.28 —— 54.51 ——
+Table 1: Percent agreement (%) and Spearman correlation ($\times$100) of different evaluation methods with human judgments on four event extraction subtasks.
+All LLM experiments are repeated three times.
+All Spearman correlation scores are associated with the $p$-value $<$ 0.05.
+For each metric in each subtask, we highlight the highest value using a dark shade, and the second-highest value using a light shade.
+3.2 Adaptive Mechanism
+Human annotators may have slightly different subjective perspectives when performing semantic evaluation, which could lead to inconsistencies in assessment Lan et al.
 (2024); Kim et al.
 (2023).
-Therefore, for both ED and EAE tasks, we manually identify several key criteria of evaluation bias, like evaluating pronouns in co-reference as correct extractions, and prompt LLMs to follow them to judge the quality of event extraction results step by step.
-More details about our human-annotated key criteria of evaluation bias and reassessment prompts can be found in Appendix C and Appendix D.1, respectively.
-It should be noted that these key criteria are optional and adaptive, contributing to the adaptive mechanism in prompts.
-This allows researchers to tailor the evaluation criteria of RAEE to their specific needs in different scenarios.
+Besides, various datasets often have customized annotation rules, resulting in differing evaluation criteria.
 For example, some datasets provide separate annotations for co-references to assist in the annotation of event extraction Li et al.
@@
 (2020).
-To achieve fair comparison in experiments, the evaluation criteria remain the same for all the evaluated models in our work.
-Task - Agree. $\uparrow$ Spearman $\uparrow$
-ED-p Human-Average 89.08 74.18
-GPT4o-Human 76.89 49.09
-      - w/o. Adaptive 78.99 47.25
-EM-Human 26.05 ——
-ED-r Human-Average 96.87 92.64
-GPT4o-Human 87.17 74.72
-      - w/o. Adaptive 84.19 69.75
-EM-Human 71.05 ——
-EAE-p Human-Average 93.61 85.47
-GPT4o-Human 88.13 72.55
-      - w/o. Adaptive 77.28 53.16
-EM-Human 63.49 ——
-EAE-r Human-Average 96.96 94.00
-GPT4o-Human 86.10 75.27
-      - w/o. Adaptive 86.42 76.08
-EM-Human 72.55 ——
-Table 1: Agreement percent and Spearman correlation with human average.
--p and -r represent the evaluation of precision and recall, respectively.
-All Spearman correlation scores are associated with the $p$-value $<$ 0.05, indicating the results are statistically significant.
-Experimental results denote that LLM achieves a moderate and strong correlation with human judgments.
-Since EM fails to achieve basic correlation scores with human judgments, their Spearman correlation scores cannot be computed correctly.
-3.3 RAEE Correlation with Human Evaluation
-To prove the effectiveness of our proposed RAEE evaluation framework and adaptive evaluation, we conduct the meta-evaluation to analyze the correlation between RAEE and human judgments.
-Specifically, we invite 3 professional event extraction annotators to evaluate 200 sampled items of data for both ED and EAE tasks on precision and recall metrics.
-These annotators are given the same prompts as those given to LLM.
-With reassessing results from 3 humans and a LLM (GPT-4o), we calculate two widely-used inter-annotator agreement (IAA) metrics among humans and LLM: percent agreement Zheng et al.
+To mitigate inconsistent evaluation and achieve adjustable evaluation in RAEE, we manually identify several key criteria of evaluation bias, and prompt the LLM to follow them to assess extractions step by step, as illustrated in Figure 2.
+When an evaluated content meets a specific evaluation criteria, the evaluation agent will explicitly output the corresponding criteria it follows.
+These key criteria are adaptive, which allows researchers to tailor the evaluation criteria of RAEE to their specific needs in different scenarios.
+In the following evaluation experiments, these key criteria remain consistent to achieve consistent evaluation.
+The default setting of existing criteria can be found in Appendix B.
+3.3 Reliability of RAEE
+To validate the reliability of RAEE and the effectiveness of its adaptive mechanism, we conduct a meta-evaluation to analyze the correlation between RAEE and human judgments.
+Specifically, we sample 800 pieces of data in ED and EAE tasks from mainstream event datasets (presented in the main experiment) across various domains, excluding extractions easily verified as correct or recalled by EM.
+We apply different LLMs as the evaluation agent to investigate the impact of agent selection on RAEE.
+The full prompts are provided in Appendix C.1.
+For human judgments, 3 professional event extraction annotators evaluate the same data using the same prompts as the LLMs.
+We also apply EM, head noun evaluation Li et al.
+(2021) and text similarity evaluation with a well-selected threshold $t=0.5$ Chen et al.
+(2023) for comparison.
+Two widely-used inter-annotator agreement (IAA) metrics are selected: Percent agreement Zheng et al.
 (2023); Liu et al.
 (2021) and Spearman correlation Lan et al.
-(2020, 2024); Fu et al.
-(2023).
-The experimental results are shown in Table 1.
-3.3.1 Correlation on Recall Task
-While LLM correlation with human average shows a noticeable drop (average 18.33 Spearman and 10.28% agreement percent), they still achieve a very strong correlation with human judgments on Agreement percent and Spearman correlation, i.e., over 86% agreement percent, which significantly outperforms EM with over 14.84% agreement percent.
-This also suggests that while the LLM can simulate human judgments and be regarded as an automatic evaluation agent, there is still room for further refinement.
-3.3.2 Correlation on Precision Task
-LLMs achieve a strong correlation with human judgments, over 76% agreement percent and 0.49 Spearman correlation of ED-p Lan et al.
 (2024); Fu et al.
 (2023).
-Furthermore, the average correlation scores between LLMs and human average for ED-p are significantly lower than that in the other tasks, indicating that semantic evaluation in this task is more challenging.
-Our investigation reveals two main reasons for this phenomenon: (1) Within the same domain, different datasets have personalized definitions for event types with similar names; (2) As for datasets across different domains, there are discrepancies in how events are defined as being triggered.
-When sampling data from multiple datasets, these subtle definitional differences can lead to confusion among evaluators when determining whether trigger predictions are accurate, leading to lower consistency in ED-p.
-In contrast, evaluating recall is simpler since it only requires determining whether gold annotations are retrieved, without judging the accuracy of the event type.
-Also, evaluating event arguments is more straightforward, as the role types for event arguments are clearly defined across datasets.
-This explains the relatively higher IAA observed in the three tasks other than ED-p.
+The results are presented in Table 1.
+3.3.1 Overall Correlation
+When using deepseek-r1 as the evaluation agent of RAEE, although there is a noticeable gap between RAEE and human average evaluation (average 0.197 Spearman correlation and 9.75% Percent agreement), RAEE still shows strong consistency with human judgments [^3] These metrics reflect a strong correlation, as validated in past works Lan et al.
+(2020); Fu et al.
+(2023)., achieving over 77% Percent agreement and 0.531 Spearman correlation.
+Compared to traditional evaluation methods, RAEE significantly outperforms exact match and head noun evaluations, improving Spearman correlation by over 0.351 across all tasks.
+Notably, the similarity-based evaluation shows little difference from RAEE on the ED recall task.
+This is because golden triggers to be evaluated in ED recall task are correct and typically single-word mentions.
+When the evaluation target shifts to model predictions that might be incorrect, as in the precision task, or involves multi-word mentions, as in the EAE task, the performance of similarity-based evaluation significantly declines.
+Furthermore, for all evaluation methods, the correlation with human judgments in precision tasks are significantly lower than those in recall tasks.
+This discrepancy arises from the inherent difficulty in evaluating precision compared to recall.
+Specifically, in the precision evaluation task, the evaluation focuses on assessing the correctness of model predictions.
+However, since the evaluated targets in the recall evaluation task are assumed to be correct, the evaluation primarily involves identifying potential correspondences.
+This phenomenon is also observed among human evaluators, as the correlation among humans is lower in the precision task compared to the recall task.
 3.3.2 Ablation of Adaptive Mechanism
-Moreover, it can also be observed that the adaptive mechanism significantly improves the percent agreement and Spearman correlation between RAEE and human average in most cases, i.e., over 2.85% percent agreement and 26.34 Spearman correlation absolute improvements.
-This is because pre-defined criteria help reduce the randomness in the LLM evaluation.
-Summary
-While LLM slightly underperforms human average, it significantly outperforms EM and achieves a strong correlation with human judgments.
-It is important to note that RAEE is a trade-off solution by considering the evaluation cost and reliability, which could reliably assess the performance of event extraction results with affordable cost.
-For example, as shown in Table 2, the average time and financial cost of GPT-4o on each evaluated model and each dataset are much smaller than those of pure human annotation, proving the scalability and reliability of LLM judge models in our proposed RAEE framework.
-Evaluation Cost Time (hours) Financial (\$)
-RAEE + GPT-4o $\approx 0.23$ $\leq 5$
-RAEE + Human $\approx 19.22$ $\approx 182.32$
-Table 2: The average evaluation cost of human annotation and GPT-4o for each evaluated model on each dataset.
+The proposed adaptive mechanism improves correlation with human in most cases, regardless of the evaluation agent.
+Besides, except for the deepseek series LLMs [^4] The technical report of deepseek mentions that its LLM is sensitive to prompts DeepSeek-AI et al.
+(2025).
+, the adaptive mechanism significantly enhances the reliability of RAEE, as evidenced by a notable reduction in standard deviations across most evaluation results.
+For example, the standard deviation for GPT-4o decreased by at least 0.36 (EAE precision) and by up to 4.46 (EAE recall) across all tasks.
+This validates the effectiveness of the adaptive mechanism.
+3.3.3 Impart of Evaluation Agents Selection
+It is observed that the evaluation quality of RAEE depends on the capability of the evaluation agent.
+Recently, deepseek-r1 has been validated as the most comprehensive model across various tasks DeepSeek-AI et al.
+(2025).
+In RAEE, deepseek-r1 delivers the overall strongest evaluation performance.
+Notably, it achieves a very strong correlation with human judgments in the EAE recall task, with a Percent agreement of 93.49% and a Spearman correlation of 0.87, significantly outperforming other LLMs.
+It is important to note that RAEE is a trade-off solution by balancing evaluation cost and reliability.
+When traditional methods fail to provide reliable semantic-level evaluation and pure human evaluation is costly, RAEE offers a relatively reliable and affordable alternative.
+As shown in Table 2, the average time and financial cost of RAEE are significantly lower than those of pure human annotation.
+Furthermore, RAEE is not restricted by the choice of specific evaluation agents.
+With the improvement in the text evaluation capabilities of LLMs, the performance of RAEE will also improve, further reinforcing its reliability.
+Evaluation Methods Time (hours) Financial (\$)
+RAEE + Deepseek-r1 $\approx 5.85$ $\approx 5.44$
+RAEE + GPT-4o $\approx 0.20$ $\approx 17.79$
+Pure Human $\approx 29.27$ $\approx 255.13$
+Table 2: The average evaluation cost of human annotation and RAEE for an evaluated model on a dataset (approximate 1170 pieces of data on average). Assume RAEE evaluates with 10 parallel programs.
 Model ACE05 Genia2011 CASIE PHEE
@@
 For each metric in each evaluation framework, we highlight the highest value using a dark shade, and the second-highest value using a light shade.
-4 Re-Assessment Experiments
-The reliability of RAEE with GPT-4o as the judging agent has been proven.
-In this section, we reassess the performance of existing event extraction models in numerous datasets across various domains by using our proposed RAEE evaluation framework.
+4 Reassessment Experiments
+The reliability of RAEE has been validated. In this section, we will apply RAEE to reassess the performance of existing EE models in mainstream datasets across various domains.
 4.1 Experimental Settings
 Tasks and Metrics
-ED and EAE event extraction tasks are evaluated respectively.
+ED and EAE tasks are evaluated separately.
 Following Lin et al.
-(2020), we focus on trigger classification (TC) and argument classification (AC).
-The semantic-level f1-score for TC and AC is described in Section 3.1, using specialized prompts to assess both precision and recall.
+(2020), we focus on trigger classification (TC) and argument classification (AC) tasks.
+The semantic-level f1-score for TC and AC is described in Section 3.1.
+Due to the time constraints, we select GPT-4o as the evaluation agent for RAEE.
 Reassessed Models
-We reassess the overall 8 state-of-the-art event extraction models with different frameworks (5 models for ED and EAE tasks).
-These models can be broadly classified into two categories: (1) Discriminative models that classify each token of text into predefined categories, including Tagprime-C Hsu et al.
+We reassess 8 state-of-the-art event extraction models using different frameworks.
+These models can be broadly categorized into two types: (1) Discriminative models that classify tokens into predefined categories, including Tagprime-C Hsu et al.
 (2023a), OneIE Lin et al.
 (2020), Query&Extract Wang et al.
-(2022), CEDAR Li et al.
-(2023a); (2) Generative models that produce event-related outputs directly, including BartGen Li et al.
+(2022), and CEDAR Li et al.
+(2023a); (2) Generative models that directly produce event-related outputs, including BartGen Li et al.
 (2021), PAIE Ma et al.
 (2022), DEGREE Hsu et al.
-(2022), AMPERE Hsu et al.
+(2022), and AMPERE Hsu et al.
 (2023b).
-All models train and inference under the unified TextEE framework Huang et al.
-(2024), with implementation details shown in Appendix E.
+Implementation details are shown in Appendix D.
 Event Datasets
-We use an overall 10 datasets to test the performance of models, covering diverse domains, including news of ACE05-en Doddington et al.
+We evaluate the performance of event models on 10 datasets covering diverse domains, including news ACE05-en Doddington et al.
 (2004) and RAMS Ebner et al.
-(2020), general domain of FewEvent Deng et al.
+(2020), general FewEvent Deng et al.
 (2020) and GENEVA Parekh et al.
-(2023), Wikipedia of MEE-en Pouran Ben Veyseh et al.
+(2023), Wikipedia MEE-en Pouran Ben Veyseh et al.
 (2022) and WikiEvents Li et al.
-(2021), biomedical domain of Genia2011 Kim et al.
-(2011), multimedia of M2E2 Li et al.
-(2020), cybersecurity of CASIE Satyapanich et al.
-(2020), and Pharmacovigilance of PHEE Sun et al.
+(2021), biomedical Genia2011 Kim et al.
+(2011), multimedia M2E2 Li et al.
+(2020), cybersecurity CASIE Satyapanich et al.
+(2020), and pharmacovigilance PHEE Sun et al.
 (2022).
-The statistics of these datasets are provided in Appendix B.
+Dataset statistics are provided in Appendix A.
 Model ACE05 Genia2011 CASIE PHEE
@@
 4.2 Main Results
-Table 3 and 4 show EM evaluations and RAEE reassessment results for ED and EAE tasks, respectively.
-Notably, across all models, RAEE reassessment shows a significant improvement in metrics compared to EM evaluations.
-On average, the precision of all models increased by more than 20%, with recall improvement exceeding 10% for ED and 15% for EAE.
-This indicates that EM evaluation severely underestimates model performance by overlooking a substantial number of semantically correct predictions, while RAEE provides a more accurate evaluation of model performance across diverse datasets.
+Table 3 and 4 show both EM evaluations and RAEE reassessment results for ED and EAE tasks, respectively.
+Notably, across all models, RAEE reassessment shows a significant improvement in metrics compared to EM.
+On average, the precision of all models increased by more than 20.52%, with recall improvement exceeding 9.85% for ED and 16.71% for EAE.
+It indicates that EM severely underestimates model performance by overlooking a substantial number of semantically correct predictions.
+Moreover, as shown in Table 5, the evaluated models’ performance gap between RAEE and EM is significantly larger for generative models compared to discriminative models, indicating that EM introduces a greater bias for generative models.
 Task Model Category Average $\Delta$ f1
@@
 Generative Models 21.47
-Table 5: Average f1-score difference between EM and RAEE evaluation. It can be found that generative models are significantly more underestimated compared to discriminative models.
-Moreover, as shown in Table 5, compared to discriminative models, the performance gap between RAEE and EM is significantly larger for generative models, indicating that EM introduces a greater bias for generative models.
-Furthermore, the rank of models across different datasets changes considerably under RAEE.
-For instance, in the ED task, DEGREE ranks lower than OneIE and CEDAR on average under EM evaluation, while DEGREE outperforms these two models under RAEE.
-This suggests that EM evaluation distorts the true performance rank of models, leading to misconceptions about the model performance.
+Table 5: Average f1-score gap between EM and RAEE.
+Furthermore, in many datasets, model rankings change considerably when shifting EM to RAEE.
+For instance, in the ED task, DEGREE ranks lower than OneIE and CEDAR across all datasets under EM, but outperforms both on average under RAEE.
+In the EAE task, rankings shift in ACE05, Genia2011 and CASIE datasets.
+It suggests that EM distorts the true performance ranks of models, leading to misconceptions about model performance.
 4.3 Fine-grained Reason Analysis
-We further conduct a fine-grained analysis to uncover the main reasons behind the misjudgments by EM, as well as the primary failure modes of existing event extraction models under our proposed RAEE evaluation framework.
-Specifically, 200 sampled data items for both ED and EAE tasks are annotated through multiple human experts.
-Reasons of EM Misjudgements
-As for instances evaluated as incorrect under EM but reassessed as correct under RAEE, we summarize 4 main categories of reasons for ED and EAE tasks, through human annotations: (1) With CoreWord: Compared to the ground-truth, the predicted results contain the core words and only have missing or redundant words that do not affect the overall semantics; (2) Co-reference: There is a relationship between the predicted result and a certain ground-truth refer to the same entity or concept; (3) Correct: While no match is found in ground-truth, these instances are still reassessed as correct; (4) Other non-representative reasons.
+To explore the main reasons behind the misjudgments evaluated by EM and the primary failure modes in RAEE, we conduct a fine-grained analysis experiment, extending the reasons reported in Han et al.
+(2024); Chen et al.
+(2024a).
+200 sampled pieces of data for both ED and EAE tasks are evaluated through multiple human experts.
+Misjudgments under EM
+As for instances evaluated as incorrect under EM but reassessed as correct under RAEE, we summarize 4 main categories of reasons: (1) With CoreWord: Compared to the ground-truth, the predicted results contain the core words and only have missing or redundant words that do not affect the overall semantics; (2) Co-reference: There is a relationship between the predicted result and a certain ground-truth refer to the same entity or concept; (3) Unannotated Correct: While no match is found in ground-truth, the predicted results are still reassessed as correct; (4) Other non-representative reasons.
 Figure 3: Distribution of the reasons that misjudges by EM evaluation method on ED and EAE tasks.
-As shown in Figure 3, it can be observed that the most misjudgments for EM evaluation in ED and EAE tasks are CoreWord and Correct.
-This indicates that current event extraction models tend to retain core words while expanding or omitting some less relevant words without changing the semantic meaning of the extracted results.
-Additionally, this also suggests that the annotations of existing datasets may not be comprehensive, as there are still a significant number of extracted results that are reasonable and correct but are not included in the ground-truth.
-Failure Modes of Existing Models Under RAEE
-As for instances evaluated as incorrect under both EM and RAEE, we also summarize 4 kinds of reasons for ED and EAE tasks, through human annotations: (1) Without CoreWord: Compared to the ground-truth, the predicted results have missing core words that convey the primary semantics; (2) WrongType: Even though there is a semantic-level match of span in the ground-truth, the predicted results have an incorrect type of event or role; (3) Incorrect: These instances are reassessed as incorrect and they are not included in the ground-truth; (4) Other non-representative reasons.
-Figure 4: Distribution of failure modes under our proposed RAEE evaluation framework on ED and EAE tasks.
-As shown in Figure 4, the most frequent failure modes of existing models on ED and EAE tasks are WrongType.
-This phenomenon indicates that although existing models could locate the spans correctly, their types are usually misclassified.This suggests that future works should focus on improving the capabilities of models in understanding and classifying the types of triggers and arguments.
+As shown in Figure 3, the main misjudgments in EM are CoreWord and U-Correct.
+It indicates that current EE models tend to preserve core words while expanding or omitting some less relevant words, without changing the semantic meaning of extractions.
+The proportion of U-Correct is similar to the 38.70% reported in Han et al.
+(2024).
+Two real examples are provided in Figure 4.
+This suggests that it may be necessary to introduce new rules, such as supplementary annotations, to enhance the comprehensiveness of some datasets.
+Figure 4: Two examples of Unannotated Correct cases.
+Failure Modes under RAEE
+As for instances evaluated as incorrect under both EM and RAEE, we also summarize 4 kinds of reasons: (1) Without CoreWord: Compared to the ground-truth, the predicted results have missing core words that convey the primary semantics; (2) WrongType: Even though there is a semantic-level match of span in the ground-truth, the predicted results have an incorrect type of event or role; (3) Unannotated Incorrect: These instances are reassessed as incorrect and they are not included in the ground-truth; (4) Other non-representative reasons.
+Figure 5: Distribution of failure modes under our proposed RAEE evaluation framework on ED and EAE.
+As shown in Figure 5, the most frequent failure mode is WrongType.
+It indicates that while existing EE models correctly locate the spans, they often misclassify the types.
+Future works should focus on enhancing models’ ability to accurately understand and classify the types of extractions.
+5 LLMs Performance in EE Task
+Previous work has validated a significant performance gap between LLMs and state-of-the-art smaller fine-tuned models in the EE task under EM Huang et al.
+(2024).
+In this section, we try to reassess LLMs performance in the EE task.
+Specifically, 6 advanced open-source and closed-source LLMs are reassessed: GPT-4o, GPT-3.5-turbo, Claude-3-Opus, Llama-3.1-70B, Qwen2-70B Yang et al.
+(2024) and Mixtral-8X7B Jiang et al.
+(2024).
+The well-selected prompts for inference are shown in Appendix C.2.
+Experiment results are shown in Table  6 and 7.
+It is found that LLMs generally underperform smaller fine-tuned models in both ED and EAE task, regardless of evaluation.
+However, LLMs show much more improvements than smaller models when shifting EM to RAEE, indicating that LLMs’ event extraction performance is greatly underestimated.
+Notably, f1-score improvement of LLMs on EAE task, over 51.92%, is higher than that on ED task, over 30.74%.
+This is mainly due to the fact that extractions in EAE task consist of more words (2.52 on average) than those in ED task (1.21 on average), making them more susceptible to misjudgment by EM.
 Model RAEE EM
@@
 Table 7: LLMs versus Top-2 state-of-the-art smaller fine-tuned models on EAE task under RAEE evaluation.
-5 Do LLMs Solve EE Task?
-Under EM evaluation system, LLMs have a significant performance gap with state-of-the-art small fine-tuned models on EE task Huang et al.
-(2024).
-In this section, we also reassess the capabilities of LLMs.
-Specifically, 6 advanced open-source and closed-source LLMs are included: GPT-4o, GPT-3.5-turbo, Claude-3-Opus, Llama-3.1-70B-Chat, Qwen2-70B-Chat Yang et al.
-(2024) and Mixtral-8X7B-instruct Jiang et al.
-(2024).
-The well-selected prompts for inference on ED and EAE tasks are shown in Appendix D.2.
-Figure 6 and 7 demonstrate results of LLMs and Top-2 fine-tuned small models under EM and RAEE evaluation on ED and EAE tasks, respectively.
-It can be found that LLMs generally underperform compared to fine-tuned models under EM evaluation, with significantly lower f1-scores on ED and EAE tasks.
-However, LLMs show remarkable improvements under RAEE evaluation, with f1-scores much closer to those of fine-tuned models, indicating that the event extraction capabilities of LLMs have been greatly underestimated.
-Notably, f1-score improvement of LLMs on EAE task, over 50%, is higher than that on ED task, over 30%.
-Our human annotations over hundreds of cases reveal that LLMs are more likely to generate out of pre-defined event type labels in ED task than role type labels in EAE task, as reflected in their relatively lower recall scores of ED.
-Additionally, we also conduct few-shot experiments with the best-performing LLM, GPT-4o [^2] https://openai.com/index/hello-gpt-4o/.
-It is found that as the number of few-shot examples increases, the performance of the model fluctuates on ED task and declines on EAE task.
+We also conduct few-shot experiments with the best-performing LLM, GPT-4o.
+It is found that as the number of few-shot examples increases, the model’s performance fluctuates on ED task and declines on EAE task.
 Our human annotations reveal that these results are heavily influenced by the quality of selected shots.
-With data to be extracted and sampled from datasets in a diverse range of domains, it is more likely to mislead the model by providing a few shots from only a few specific domains.
-We will leave a more detailed analysis of few-shot prompting strategies in LLMs for our future work.
+Since the data to be extracted is sampled from datasets across diverse domains, using few-shot examples from only a limited set of domains is more likely to mislead the model.
 6 Conclusion and Future Works
-In this paper, we propose a reliable automatic evaluation framework, RAEE, for assessing the event extraction task at the semantic level rather than the token level of exact match evaluation.
-Extensive experimental results demonstrate that GPT-4o achieves a strong correlation with human judgments within our RAEE framework, establishing RAEE as a reliable automatic evaluation framework for assessing the quality of event extraction results.
-Additionally, we reassess the performance of numerous prior works and advanced LLMs in the event extraction task using our proposed RAEE framework.
-Extensive experimental results indicate that the capabilities of prior works and LLMs have been significantly underestimated.
-Finally, the fine-grained reason analysis reveals several intriguing phenomena worth further exploration.
-In the future, we will conduct a more comprehensive evaluation for event extraction tasks to reassess the performance of existing models.
-Besides, we will extend the evaluation of model performance with RAEE from closed domains to the open domain.
-Appendix A Limitations
-A.1 LLM Judgments v.s. Human Judgments
-In this paper, advanced LLM GPT-4 is used to automatically judge the correctness of the predicted trigger words and arguments for the event extraction task.
+In this paper, we introduce RAEE, a reliable semantic-level evaluation framework for event extraction, which moves beyond the token-level exact match evaluation.
+Extensive experiments validate RAEE’s strong correlation with human judgments, confirming its reliability.
+We apply RAEE to reassess the performance of numerous prior works and advanced LLMs in EE.
+Extensive experiments indicate that both prior works and LLMs are significantly underestimated.
+A further fine-grained reason analysis reveals several intriguing phenomena worth further exploration.
+In the future, we will extend the evaluation of model performance with RAEE from closed domains to the open domain.
+Limitations
+LLM Judgments v.s. Human Judgments
+In this paper, the advanced LLM GPT-4o is used to automatically judge predictions from event extraction models.
 However, it should be noted that LLMs cannot always perfectly align with human judgments, which is the most reliable evaluation.
-For example, GPT-4o achieves comparable agreement and correlations with human average performance in ED and EAE annotations.
-Since the ED precision task is challenging, the Spearman correlation score on it is significantly lower than human performance.
+For example, GPT-4o achieves comparable agreement and correlations with human average performance in ED and EAE evaluations.
 While we have made efforts to improve the relevance between LLM and human assessment through prompt design, this gap still exists.
@@
 (2023).
-A.2 Computational Cost of Evaluating One Model
-Unlike existing EM evaluation, our proposed metrics rely on advanced LLMs to annotate the quality of predicted events, leading to much higher costs.
-The average cost of evaluating one model on each benchmark is between 2-5\$, which is smaller than established benchmarks like AlpacaEval Li et al.
+Sensitivity of LLMs to Prompts
+In our preliminary study, it is found that the evaluation agent LLMs are sensitive to prompts.
+In the meta-evaluation experiments, using different prompts significantly affects LLMs’ correlation with human judgments.
+To mitigate this problem, we provide well-designed prompts that performs well across all LLMs in the meta-evaluation experiments, which can be found in Appendix C.1.
+Computational Cost of Evaluation
+Unlike existing EM evaluation, our proposed RAEE relies on advanced LLMs to evaluate the quality of predicted events, leading to much higher costs.
+The average cost of evaluating one model on a dataset is between 2-5\$, which is smaller than established benchmarks like AlpacaEval Li et al.
 (2023b).
-A.3 Insufficient Evaluation for LLMs
+Insufficient Evaluation for LLMs
 The number of the test set in existing event extract datasets contains thousands of samples, leading to a huge cost for LLMs to do inference experiments.
@@
 Therefore, the conclusions of our experimental results are reliable and robust.
-A.4 Other Advanced LLMs as Evaluator
-In this paper, the GPT-4 model is used to conduct the automatic evaluation in RAEE.
-Due to the budget limitation, we did not employ a diverse set of advanced LLMs for this task, such as Claude-3 and PaLM.
-Since our preliminary study reveals that GPT-4o is slightly better than other advanced LLMs, we mainly use it as the automatic judge model in our proposed RAEE framework.
-Appendix B Datasets Statistics
+Appendix A Datasets Statistics
 We utilize 10 widely-used datasets as evaluation benchmarks in both EM and RAEE evaluation.
-Since these datasets come from different domains and vary in pre-defined ontologies and size, it is necessary to introduce basic statistics of these datasets, as shown in Table 8.
+Since these datasets originate from various domains and differ in terms of pre-defined ontologies as well as size, it is necessary to introduce basic statistics of these datasets, as shown in Table 8.
 Dataset #Inst #ET #Evt #RT #Arg
@@
 Table 8: Statistics of 10 datasets used in RAEE. #Inst, #ET, #Evt, #RT, and #Arg represent the number of data instances, event type, event instances, role types, and argument instances, respectively.
-Appendix C Key Criteria in RAEE
+Appendix B Key Criteria in RAEE
 Considering that the requirements for EE vary across different scenarios and applications, we introduce key criteria in RAEE to enable the adaptive mechanism.
 As shown in Tabel 9, these key criteria are based on factors that significantly influence the evaluation results in our human annotation meta-evaluation experiment.
-Furthermore, this adaptive mechanism is easily extendable, allowing for the addition of personalized judgment criteria to account for more factors in extraction.
+Furthermore, this adaptive mechanism is easily extendable, allowing for the addition of specific judgment criteria to account for more factors in extraction.
 Key Criteria Template Setting
@@
 Table 9: Key criteria templates and their settings used in RAEE.
-Appendix D Prompts
-D.1 Reassessment Prompts in RAEE
+Appendix C Prompts
+C.1 Reassessment Prompts in RAEE
 Table 10, 11, 13 and 15 illustrate the prompts we use for reassessment in RAEE.
-D.2 Inference Prompts for LLMs Extracting Events
+C.2 Inference Prompts for LLMs Extracting Events
 Table 12 and 14 illustrate the prompts we use for testing the event extraction ability of LLMs in ED and EAE tasks, respectively.
+It should be noted that we follow existing closed domain EE studies Gao et al.
+(2023); Wei et al.
+(2024); cai-etal-2024-improving-event to explicitly provide all possible types in the inference prompt.
+This is a necessary step when performing EE inference with LLMs and does not constitute information leakage.
 Prompt Used for Reassessing on ED Task and Precision Metric
@@
 Table 14: Prompt used for testing the event extraction ability of LLMs in EAE task. An example of data is illustrated.
-Appendix E Implementation Details
+Appendix D Implementation Details
 To avoid inconsistency in evaluation caused by factors out of model frameworks such as data preprocessing or Python package versions, we follow TextEE Huang et al.